OpenStudy (anonymous):

5y-4=9y+8 Solve equation, show work please

OpenStudy (gw2011):

5y-4=9y+8 -4-8=9y-5y -12=4y Rewriting this you get: 4y=-12 y=-12/4 y=-3

OpenStudy (anonymous):

the sum of six and four times d is the same as d minus nine. translate each sentence into an equation

OpenStudy (anonymous):

help me